Odds Ratio Explained – Clinical Applications in Psychiatry & Research

Posted on: Last Updated:
2-4 minutes 4,930
Download PDF

Odds

Odds of an event happening is defined as the likelihood that an event will occur, expressed as a proportion of the likelihood that the event will not occur. Therefore, if A is the probability of subjects affected and B is the probability of subjects not affected, then odds = A /B.

Therefore, the odds of rolling four on dice are 1/5 or an implied probability of 20%. [Note this is not the same as probability which would be 1/6 = 16.66%]

Odds Ratio (OR) is a measure of association between exposure and an outcome. The OR represents the odds that an outcome will occur given a particular exposure, compared to the odds of the outcome occurring in the absence of that exposure.

Important points about Odds ratio:

  • Calculated in case-control studies as the incidence of outcome is not known
  • OR >1 indicates increased occurrence of an event
  • OR <1 indicates decreased occurrence of an event (protective exposure)
  • Look at CI and P-value for statistical significance of value (Learn more about p values and confidence intervals here)
  • In rare outcomes OR = RR   (RR = Relative Risk). This applies when the incidence of the disease is < 10%

If you are interested in doing a full research and statistics course for critical analysis/ appraisal then click here. 

Also, don’t forget to download our free critical appraisal worksheets to analyse research papers.

Enhance Your Psychiatry Practice

Join The Academy by Psych Scene and earn CME & CPD points from 150+ hours of expert-led psychiatry courses. Hub Pro Included ✅

Check out our courses
Blue box image

CLINICAL EXAMPLE AND CALCULATION

In a study examining the association between estrogen (exposure) and endometrial carcinoma (outcome), the two by two table is shown below.

Learning point: In a two by two table, for ease of calculation ensure that the outcome of interest is always at the top and the exposure on the left.

Odds ratio calculation

Interpretation

According to the tablet above, individuals with endometrial cancer are 4.42 times more likely to be exposed to estrogen than those without endometrial carcinoma.

Learning point: It is not appropriate to interpret this as ‘Individuals with estrogen exposure are 4.42 times more likely to develop Endometrial cancer than those without exposure.’ The reason is that a case-control study begins from outcome i.e. selection of a sample with the outcome of interest which in this case is endometrial cancer.

Get Serious About Psychiatry Learning

Our courses offer practical knowledge and clinical expertise at exceptional value, plus CME & CPD points.

Go to Academy by Psych Scene
Blue box image